# Impact of Dataset Properties on Membership Inference Vulnerability of Deep Transfer Learning

### v5

@@ A.6 Proof of Theorem˜2
$\mu_{\mathrm{in}}-\mu_{\mathrm{out}}$ $=\mathbb{E}[s_{y}^{\mathrm{(in)}}-s_{y}^{\mathrm{(out)}}]=\frac{1}{S}\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle$ (A55)
@@
$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})$ $=\Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))+\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right),$ (A57)
@@
$r=\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}.$ (A58)
@@
$=\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})+r\phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}r^{k}.$ (A63)
It follows that
@@
$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}$ (A66)
$+\log\left(\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}\left(\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{k}\right)$ (A67)
$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}$ (A68)
$+\log\left(1+\sum_{k=1}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}\left(\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{k}\right)$ (A69)
$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S)),$ (A70)
@@
$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (A71)
@@ A.7 Proof of Corollary˜4
$\approx\int_{\mathscr{D}}p(\bm{x})\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\mathrm{d}\bm{x}$ (A73)
$=\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\frac{1}{\sqrt{S}}\int_{\mathscr{D}}p(\bm{x})\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\mathrm{d}\bm{x}$ (A74)
$=\frac{1}{\sqrt{S}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{2\pi}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right],$ (A75)
@@
$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}+\log\left(\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{2\pi}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right]\right).$ (A76)
@@ B.3 RMIA power-law
To apply Lemma˜8, we will calculate $\mathbb{E}_{\bm{z}}[|q|]$ and $\mathbb{E}_{\bm{z}}[|A|]$. Let $\mathcal{D}_{\mathrm{target}}=\{(\bm{x}_{j,1},j),\dots,(\bm{x}_{j,S},j)\}_{j=1}^{C}$. Let $s_{y_{a}}^{(\bm{b})}(\bm{a})$ denote the score of $(\bm{a},y_{\bm{a}})$ when $\mathcal{D}_{\mathrm{target}}$ contains $(\bm{b},y_{\bm{b}})$ but not the other example. Using similar argument as in the proof of Theorem˜2, we have
@@
$\mu_{\bm{x},\bm{x}}$ $=\langle\bm{x},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{x},\bm{x}-m_{y_{\bm{x}}}\rangle$ (A131)
@@
$\mu_{\bm{x},\bm{z}}$ $=\begin{cases}\langle\bm{z},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{z},\bm{x}-m_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \langle\bm{z},m_{y_{\bm{z}}}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A133)
$\mu_{\bm{z},\bm{z}}$ $=\langle\bm{z},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{z},\bm{z}-m_{y_{\bm{x}}}\rangle\quad$ (A134)
@@
where $m_{y_{\bm{z}}}$ is the true class mean of class $y_{\bm{z}}$.
@@
$+\frac{C-1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left(\frac{\langle\bm{x},\bm{x}-m_{y_{\bm{x}}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{2}+\left(\frac{\langle\bm{z},\bm{z}-m_{y_{\bm{z}}}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right)^{2}\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right]$ (A143)
@@
$+\frac{C-1}{CS}\mathbb{E}_{\bm{z}}\left[\left.\frac{\langle\bm{x},\bm{x}-m_{y_{\bm{x}}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-m_{y_{\bm{z}}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right],$ (A145)
@@
$+\frac{C-1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right]$ (A148)
@@
$+\frac{C-1}{C\sqrt{S}}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right].$ (A150)
@@
$\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\\ =\frac{1}{\sqrt{S}}\cdot\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A151)
@@
$\Psi(\bm{x},C)=\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A154)
@@
As for the LiRA power-law, bounding $\bm{x}-m_{\bm{x}}||$ and $\bm{z}-m_{\bm{z}}||$ will provide a worst-case upper bound for which the power-law holds.

### v6

@@ A.6 Proof of Theorem˜2
$\mu_{\mathrm{in}}-\mu_{\mathrm{out}}$ $=\mathbb{E}[s_{y}^{\mathrm{(in)}}-s_{y}^{\mathrm{(out)}}]=\frac{1}{S}\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle$ (A55)
@@
$\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})$ $=\Phi\left(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))+\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right),$ (A57)
@@
$r=\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}.$ (A58)
@@
$=\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})+r\phi(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}r^{k},$ (A63)
where $\mathrm{He}$ denotes Hermite polynomials. It follows that
@@
$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}$ (A66)
$+\log\left(\sum_{k=0}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}\left(\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{k}\right)$ (A67)
$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}$ (A68)
$+\log\left(1+\sum_{k=1}^{\infty}\frac{(-1)^{k}\mathrm{He}_{k}(\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x})))}{(k+1)!}\left(\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{k}\right)$ (A69)
$=$ $-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S)),$ (A70)
@@
$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (A71)
@@ A.7 Proof of Corollary˜4
$\approx\int_{\mathscr{D}}p(\bm{x})\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\mathrm{d}\bm{x}$ (A73)
$=\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\frac{1}{\sqrt{S}}\int_{\mathscr{D}}p(\bm{x})\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\mathrm{d}\bm{x}$ (A74)
$=\frac{1}{\sqrt{S}}e^{-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}}\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{2\pi}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right],$ (A75)
@@
$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}+\log\left(\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{2\pi}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right]\right).$ (A76)
@@ B.3 RMIA power-law
To apply Lemma˜8, we will calculate $\mathbb{E}_{\bm{z}}[|q|]$ and $\mathbb{E}_{\bm{z}}[|A|]$. Let $\mathcal{D}_{\mathrm{target}}=\{(\bm{x}_{j,1},j),\dots,(\bm{x}_{j,S},j)\}_{j=1}^{C}$. Let $s_{y_{\bm{a}}}^{(\bm{b})}(\bm{a})$ denote the score of $(\bm{a},y_{\bm{a}})$ when $\mathcal{D}_{\mathrm{target}}$ contains $(\bm{b},y_{\bm{b}})$ but not the other example. Using similar argument as in the proof of Theorem˜2, we have
@@
$\mu_{\bm{x},\bm{x}}$ $=\langle\bm{x},\bm{m}_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{x},\bm{x}-\bm{m}_{y_{\bm{x}}}\rangle$ (A131)
@@
$\mu_{\bm{x},\bm{z}}$ $=\begin{cases}\langle\bm{z},\bm{m}_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{z},\bm{x}-\bm{m}_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \langle\bm{z},\bm{m}_{y_{\bm{z}}}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A133)
$\mu_{\bm{z},\bm{z}}$ $=\langle\bm{z},\bm{m}_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{z},\bm{z}-\bm{m}_{y_{\bm{x}}}\rangle\quad$ (A134)
@@
where $\bm{m}_{y_{\bm{z}}}$ is the true class mean of class $y_{\bm{z}}$.
@@
$+\frac{C-1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left(\frac{\langle\bm{x},\bm{x}-\bm{m}_{y_{\bm{x}}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{2}+\left(\frac{\langle\bm{z},\bm{z}-\bm{m}_{y_{\bm{z}}}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right)^{2}\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right]$ (A143)
@@
$+\frac{C-1}{CS}\mathbb{E}_{\bm{z}}\left[\left.\frac{\langle\bm{x},\bm{x}-\bm{m}_{y_{\bm{x}}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-\bm{m}_{y_{\bm{z}}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right],$ (A145)
@@
$+\frac{C-1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-\bm{m}_{\bm{z}}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right]$ (A148)
@@
$+\frac{C-1}{C\sqrt{S}}\mathbb{E}_{\bm{z}}\left[\left.\left|\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-\bm{m}_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\ \right|\ y_{\bm{z}}\not=y_{\bm{x}}\right].$ (A150)
@@
$\frac{\mathbb{E}_{\bm{z}}[|q|]}{\mathbb{E}_{\bm{z}}[|A|]}\\ =\frac{1}{\sqrt{S}}\cdot\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-\bm{m}_{\bm{z}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-\bm{m}_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A151)
@@
$\Psi(\bm{x},C)=\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle^{2}}{\bm{x}^{T}\Sigma\bm{x}}+\frac{\langle\bm{z},\bm{z}-\bm{m}_{\bm{z}}\rangle^{2}}{\bm{z}^{T}\Sigma\bm{z}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}}+\frac{\langle\bm{z},\bm{z}-\bm{m}_{\bm{z}}\rangle}{\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A154)
@@
As for the LiRA power-law, bounding $\bm{x}-\bm{m}_{\bm{x}}||$ and $\bm{z}-\bm{m}_{\bm{z}}||$ will provide a worst-case upper bound for which the power-law holds.